How would I Factor this equation,if f(t)=2t^2-4t-1,find- with f(2) and f(1).
Do, I replace (t), with 2 and 1 in the equation. And, thanks for the help.

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

yes __ just substitute the values of t to find the value of the function

Answer by stanbon(75887)

How would I Factor this equation,if f(t)=2t^2-4t-1,find- with f(2) and f(1).
Do, I replace (t), with 2 and 1 in the equation.
-----------------
2t^2-4t-1
Use the quadratic formula:
t = [4 +- sqrt(16 - 4*2*-1)]/4
t =[4 +- sqrt(24)]/4
t =[4 +- 2sqrt(6)]/4
t = 1 + (1/2)sqrt6 or t = 1 - (1/2)sqrt6
Factors:
(t-1-(1/2)sqrt6)(t-1+(1/2)sqrt6)
---------------
f(t)=2t^2-4t-1,find- with f(2) and f(1).
Do, I replace (t), with 2 and 1 in the equation.
Yes, that is what you do.

=====================
Cheers,
Stan H.